Box A and Box B both contain the numbers 1,2,3, and 4. Construct the probability mass function and draw the histogram of the sum when one number from each box is taken at a time, with replacement


1
Expert's answer
2022-03-17T18:48:07-0400

Let X be a random variable representing the sum of the numbers, then X = {2,3,4,5,6,7,8}

"P(X=2)={\frac 1 {16}}"

"P(X=3)={\frac 2 {16}}={\frac 1 8}"

"P(X=4)={\frac 3 {16}}"

"P(X=5)={\frac 4 {16}}={\frac 1 4}"

"P(X=6)={\frac 3 {16}}"

"P(X=7)={\frac 2 {16}}={\frac 1 8}"

"P(X=8)={\frac 1 {16}}"